an elevation car whose floor to ceiling distance is 2.7 metre starts ascending with constant acceleration of 1.2 m/s^2. 2 sec after start a ball begins to fall from the ceiling of the car.find the free fall time of the ball

Grade:11

1 Answers

Kurma Anudeep
11 years ago

t=2s

d=2.7m

a=1.2m/s^2

a=dv/dt

1.2=2.7v/5.4

v=2.4m/s

v=ds/dt

2.4=d(2.7)/dt

t=2.7/2.4

t=1.125s